### v1

# Scaling Vision Mamba Across Resolutions via Fractal Traversal

Abstract
Vision Mamba has recently emerged as a promising alternative to Transformer-based architectures, offering linear complexity in sequence length while maintaining strong modeling capacity. However, its adaptation to visual inputs is hindered by challenges in 2D-to-1D patch serialization and weak scalability across input resolutions. Existing serialization strategies such as raster scanning disrupt local spatial continuity and limit the model’s ability to generalize across scales. In this paper, we propose FractalMamba++, a robust vision backbone that leverages fractal-based patch serialization via Hilbert curves to preserve spatial locality and enable seamless resolution adaptability. To address long-range dependency fading in high-resolution inputs, we further introduce a Cross-State Routing (CSR) mechanism that enhances global context propagation through selective state reuse. Additionally, we propose a Positional-Relation Capture (PRC) module to recover local adjacency disrupted by curve inflection points. Extensive experiments on image classification, semantic segmentation, object detection, and change detection demonstrate that FractalMamba++ consistently outperforms previous Mamba-based backbones, particularly under high-resolution settings.
Index Terms:
Vision Mamba, Fractal Scanning, Cross-State Routing, Positional-Relation Capture
1 Introduction
The field of artificial intelligence has recently seen a significant shift towards the development and deployment of foundation models [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. These models, characterized by their ability to learn generalizable representations through extensive pre-training and adapt to diverse downstream tasks, have become central to progress across numerous domains [16, 17, 18, 19]. At the core of foundation models lies the Transformer architecture [20]. Despite their remarkable success, Transformers face a critical limitation stemming from the self-attention mechanism: its computational and memory complexity scales quadratically with the input sequence length. This quadratic scaling poses a significant bottleneck, rendering Transformers computationally expensive and memory-intensive when processing long sequences, which motivates the search for fundamentally different architectures.
Recently, Mamba has emerged as a promising alternative, offering a way to model long sequences with greater efficiency [21, 22]. Crucially, Mamba maintains linear time complexity in sequence length for both training and inference, overcoming the efficiency limitations of Transformers. This architecture has been adapted for computer vision in recent models [23, 24, 25, 26, 27], which typically employ linear scanning of image patch sequences to handle the non-causal nature of visual data.
However, adapting vision Mamba to the inherently non-causal, 2D structure of images presents significant challenges. A primary issue is the need to convert the 2D grid of image patches into a 1D sequence suitable for the Mamba’s input. Standard serialization methods used in existing models [23, 24, 25, 26, 27], such as raster scanning (row-by-row or column-by-column), often disrupt the crucial spatial locality present in images. Patches that are close neighbors in the 2D image grid can become distant in the 1D sequence, hindering the model’s ability to effectively learn local patterns and relationships, which are fundamental to visual understanding.
Furthermore, similar to vision Transformers (ViTs) [28, 29, 30], effectively handling inputs of varying resolutions remains a critical challenge for vision Mamba. The need for robust multi-scaling adaptability, the ability to handle varying image resolutions effectively, is paramount for practical computer vision applications where image sizes naturally vary [31]. Standard ViTs often exhibit performance degradation when encountering resolutions different from their fixed pre-training resolution. Vision Mamba, despite its linear complexity advantage, likely faces analogous difficulties. By flattening 2D patches into a 1D sequence, vision Mamba’s ability to interpret spatial relationships may be sensitive to changes in the underlying grid structure caused by varying resolutions.
Figure 1: Visualization of correlation between the final state and intermediate state tokens in the deepest feature layer under different input resolutions. The 16×16$\times$16 × and 32×32$\times$32 × maps correspond to the output token grids for input resolutions of $512\times 512$ and $1024\times 1024$, respectively. Without CSR, correlations are biased and localized due to information fading across long sequences. With CSR enabled, strong and distributed correlations are observed, demonstrating that global contextual information is effectively propagated across the sequence. “W/o” means without operation.
To address above limitations, we propose a novel approach that enhances vision Mamba’s adaptability and robustness in this paper. Firstly, to better handle the 2D-to-1D conversion while respecting the non-causal nature and spatial structure of images, we move away from simple linear scanning methods. Inspired by our previous work FractalMamba published in AAAI2025 [32], we leverage fractal space-filling curves, such as the Hilbert curve, for image patch serialization. These curves are specifically designed to traverse a multi-dimensional grid while maximizing locality preservation. By ensuring that spatially adjacent image patches remain close in the 1D sequence, fractal scanning provides the sequence model with an input that more faithfully represents the original 2D spatial relationships. This is crucial for models like vision Mamba, whose state transitions operate sequentially, allowing them to better capture local visual context. Additionally, the inherent self-similar nature of fractal curves offers superior multi-scale adaptability. They can adapt seamlessly to varying image resolutions while preserving their locality properties, unlike fixed raster scans which break down as dimensions change.
While fractal scanning improves the spatial coherence of the input image, the underlying sequential nature of the Mamba architecture presents another challenge. In state space models (SSMs), each output is computed through a chain of recurrent state transitions, where the current state is influenced primarily by recent inputs. As the sequence length increases, particularly in high-resolution images, information from earlier patches may gradually fade, limiting the model’s ability to capture long-range dependencies and global context. The first column of Fig. 1 shows an underlying heatmap encoding the correlation between the SSM’s final hidden state and each intermediate state. From this heatmap, one can observe that when image patch blocks are connected solely by the standard fractal curve, strong correlations remain confined to states adjacent along the curve, and long-range dependencies vanish. To address this issue, we introduce a Cross-State Routing (CSR) mechanism that enhances long-range information propagation. By selectively routing earlier hidden states to later stages in the sequence, this mechanism reinforces the flow of global context and helps the model retain visual cues from distant positions, as shown in the second column of Fig. 1.
Figure 2: Illustration of the structural limitation of fractal curves. Certain corner points along the fractal traversal path (highlighted by purple circles) experience disrupted adjacency relationships. Although spatially close in the 2D image grid, these patches become distant in the serialized sequence, impairing local continuity and information exchange.
Figure 3: Top-1 classification accuracy of Mamba-based models across different input resolutions on ImageNet-1K. Results are grouped by parameter scale: $\sim$30M, $\sim$11M, and $\sim$7M. Each curve represents model performance from $224^{2}$ to $1536^{2}$ input resolution. FractalMamba++ consistently outperforms other methods, especially under large image sizes, demonstrating strong resolution scalability. Comparison methods contain VMamba [26], GrootV [27], MILA [33], MSVMamba [34], SpatialMamba [35], MambaVision [36], Mamba® [37], EfficientVMamba [38], ViM [25] and FractalMamba [32].
Fractal curves, such as the Hilbert curve, preserve the local 2D structure when mapping image patches into a 1D sequence, making them highly effective for maintaining spatial continuity during serialization. However, despite their strong locality-preserving properties, fractal curves can still exhibit minor locality disruptions at corner points. As illustrated in Fig. 2, patches located at purple and red areas, which are spatially adjacent in the 2D space, may become separated in the serialized order. To mitigate this issue, our previous FractalMamba introduces a shifting mechanism that applies multiple offset scan paths, effectively bringing certain corner-adjacent patches closer together. Nevertheless, while this approach partially alleviates locality breaks, it incurs additional complexity by requiring multiple scanning curves. Recognizing that perfect 2D-to-1D mapping cannot be achieved by any single scanning curve alone, we propose to further enhance the structural modeling by embedding explicit positional-relation capture (PRC) into the patch embeddings. The PRC module serves as a complementary mechanism to reinforce local adjacency, allowing the network to recover fine-grained spatial relationships without relying solely on scan-based solutions. As shown in Fig. 3, our proposed FractalMamba++ achieves superior adaptability to input images of varying resolutions compared to existing models. The contributions of this paper are listed as follows:
  • We propose a fractal-based patch serialization strategy using Hilbert curves to convert 2D image grids into 1D sequences, effectively preserving spatial locality and improving the model’s ability to capture local visual structures. We demonstrate that the self-similar and scale-invariant properties of fractal curves enable Vision Mamba to maintain strong performance across varying input resolutions, enhancing its multi-resolution adaptability.
  • We introduce a unified enhancement that includes CSR mechanism for long-range information propagation and PRC mechanism to mitigate locality disruption caused by curve inflection, thereby further improving the model’s capacity to capture both global context and spatial adjacency.
  • We evaluate the efficacy of FractalMamba++ across various-resolutions vision tasks: image classification task, semantic segmentation task, remote sensing detection task and common object detection task. The experimental results unequivocally demonstrate enhanced performance and broad applicability of our proposed FractalMamba++.
The main differences between the FractalMamba and the FractalMamba++ are reflected in following aspects:
  • We introduce the Cross-State Routing (CSR) module to enhance global modeling by alleviating the information decay issue in long sequences, significantly improving long-range dependency learning in high-resolution images.
  • We design the Positional-Relation Capture (PRC) module to compensate for locality disruption at fractal curve inflection points, improving local detail awareness without increasing curve complexity.
  • We conduct a more comprehensive set of experiments than prior work, evaluating FractalMamba++ across multiple parameter scales on diverse visual tasks including classification, segmentation, detection, and change detection.
2 Related Work
2.1 Vision Backbone Architecture
In the domain of vision backbone architectures, significant advancements have been achieved through the development and refinement of several key frameworks. The two most widely used backbone families are CNN-based and ViT-based models, each offering distinct advantages across different computer vision tasks. CNNs [39, 40, 41, 42, 43] have historically formed the foundation of visual modeling due to their inductive biases toward local connectivity and spatial hierarchies. Classic architectures such as AlexNet [39], VGG [40], and ResNet [41] establish performance benchmarks that continue to influence model design today. More recently, ViTs [20, 30, 28, 44] introduce a paradigm shift by applying the self-attention mechanism to non-overlapping image patches, treating them as tokenized sequences similar to NLP models. ViTs are capable of modeling global context and long-range dependencies, and have become a foundational component in large-scale pretraining frameworks [1, 2, 3, 4, 5, 6, 7, 45, 9].
Despite their representational power, ViTs face notable limitations in efficiency due to their quadratic complexity with respect to sequence length, which hinders scalability to higher resolutions or longer sequences. Furthermore, ViTs often struggle to generalize across input resolutions, as their positional embeddings are learned for fixed patch grids and do not naturally adapt to scale changes [31]. These issues motivate the exploration of more resolution-flexible and computationally efficient vision backbones. Our work, FractalMamba++, contributes to this evolving landscape by proposing a vision backbone that combines the efficiency of Mamba with fractal-based spatial structure modeling and resolution adaptability. By rethinking patch serialization and positional representation, our approach offers a new perspective on building scalable, locality-aware, and resolution-robust vision backbones beyond traditional CNNs and Transformers.
Figure 4: The architecture of our FractalMamba++ Backbone. It mainly contains positional-relation capture (PRC), Fractal Serialization and Cross-State Routing (CSR) mechanism.
2.2 State Space Model
SSMs have recently emerged as promising alternatives to Transformers in sequence modeling, offering linear computational complexity in sequence length. Models such as S4 [46] and Mamba [21] replace explicit attention mechanisms with parameterized state transitions, enabling efficient long-range dependency modeling. In particular, Mamba introduces input-dependent selective updates to improve both expressiveness and scalability. Following their success in language modeling, SSMs have been extended to vision tasks [26, 27, 33, 34, 35, 36, 37, 38, 25, 32], where images are typically divided into non-overlapping patches and serialized into sequences. Similarly, SSMs are gaining momentum in image generation and vision-language tasks [47, 48], where efficient sequence modeling becomes critical for scaling to larger resolutions and longer contexts. As modern generative frameworks seek to synthesize complex visual content such as images, videos, and 3D structures, they demand architectures that can model spatial and temporal dependencies over extended input lengths. Transformers, while powerful, face growing computational challenges as resolution and sequence length increase due to their quadratic attention complexity. In contrast, Mamba demonstrate strong potential in modeling dense and long-range visual sequences with improved efficiency. This makes vision Mambas increasingly promising for tasks such as high-resolution image generation tasks, where scalable context modeling is critical.
However, unlike natural language, visual data is inherently two-dimensional. Naive rasterization methods used to serialize image patches into sequences often break spatial adjacency and distort structural continuity. This challenge affects both recognition and generation tasks, particularly under high-resolution settings, where maintaining spatial coherence is crucial for accurate modeling. Therefore, an effective serialization strategy that respects image structure becomes essential for fully exploiting the potential of SSMs in vision domains. To address this challenge, we propose FractalMamba++, a vision Mamba framework that introduces the fractal-based serialization paradigm. Instead of conventional linear scans, we leverage self-similar fractal space-filling curves such as the Hilbert curve to convert 2D images into 1D sequences while preserving spatial locality. This design ensures that adjacent patches in the image remain close in the serialized sequence, thereby facilitating more accurate spatial reasoning. Moreover, we extend the standard SSM architecture with two additional modules: a Cross-State Routing (CSR) mechanism that improves long-range dependency modeling by explicitly propagating earlier hidden states, and a Positional-Relation Capture (PRC) module that reinforces local structural awareness in patch embeddings. Together, these innovations form the backbone of FractalMamba++, which demonstrates robust adaptability to varying resolutions and tasks. In summary, FractalMamba++ builds upon state space models by incorporating fractal-based serialization, long-range state propagation, and structure-aware embedding design. These components jointly enhance the model’s ability to preserve locality, maintain global context, and adapt to variable input resolutions. The architecture of FractalMamba++ is shown in Fig. 4.
Figure 5: Visualization of patch-wise SDS for linear and fractal scanning curves. Darker colors represent higher SDS values, suggesting more severe disruption of spatial structure.
TABLE I: The percentage of patches under SDS threshold across scanning methods. All patches under the fractal curve maintain low SDS values, highlighting its strong structural consistency advantage.
SDS 1.1 1.2 1.3 1.4 1.5 1.6 1.9 2.1 2.4 2.5 4.0
Linear Scan 0% 0% 0% 3.1% 3.1% 56.2% 56.2% 56.2% 56.2% 56.2% 78.1%
Local Scan 0% 0% 0% 3.1% 3.1% 6.2% 43.8% 50% 87.5% 93.8% 96.9%
Fractal Scan 0% 3.1% 62.5% 62.5% 100% 100% 100% 100% 100% 100% 100%
3 Method
In this paper, we propose FractalMamba++, a resolution-scalable vision backbone based on state space models. Our design addresses the key challenge of preserving spatial structure and global context when adapting Mamba to two-dimensional visual inputs. Specifically, FractalMamba++ enhances vision Mamba in three aspects: spatial locality preservation, long-range dependency modeling, and structural relation encoding.
First, we adopt a Hilbert fractal curve to serialize 2D image patches into a 1D sequence. Compared to traditional raster scanning, the Hilbert curve exhibits superior locality-preserving properties, ensuring that adjacent patches in the 2D space remain close in the serialized sequence. This design allows the model to better retain spatial continuity, which is essential for maintaining consistent performance across varying image resolutions. Second, to address the inherent limitation of sequential recurrence in Mamba, where early patch information may fade over long sequences, we introduce a Cross-State Routing (CSR) mechanism. CSR enables selective propagation of earlier hidden states to later positions, thereby improving the model’s ability to capture long-range dependencies and maintain global semantic consistency. Finally, while fractal scanning preserves local proximity, it does not explicitly encode complex spatial relationships such as structural symmetry or corner adjacency. To this end, we propose a Positional-Relation Capture (PRC) module that injects spatial relational priors directly into the patch embeddings. This complementary mechanism allows the model to learn fine-grained geometric patterns and further enhances its structural awareness.
Together, these components form a unified framework that improves resolution robustness, spatial structure preservation, and long-range visual understanding in state space vision models.
3.1 Preliminaries
State Space Models (SSMs). SSMs offer a principled approach to modeling dynamical systems, particularly those governed by Linear Time-Invariant (LTI) dynamics. The continuous-time formulation of an LTI system is given by:
$\begin{split}h^{\prime}(t)&=\mathbf{A}h(t)+\mathbf{B}x(t),\\ y(t)&=\mathbf{C}h(t),\end{split}$ (1)
where $h(t)\in\mathbb{R}^{N}$ is the hidden state at time $t$, $x(t)\in\mathbb{R}$ is the input, and $y(t)\in\mathbb{R}$ is the output. The matrices $\mathbf{A}\in\mathbb{R}^{N\times N}$, $\mathbf{B}\in\mathbb{R}^{N\times 1}$, and $\mathbf{C}\in\mathbb{R}^{1\times N}$ define the system dynamics and projection functions.
Since real-world data such as images and text are inherently discrete, applying SSMs to these domains requires discretizing the continuous formulation. A standard discretization method is zero-order hold (ZOH), which transforms $\mathbf{A},\mathbf{B}$ into:
$\begin{split}\overline{\mathbf{A}}&=\exp(\Delta\mathbf{A}),\\ \overline{\mathbf{B}}&=(\Delta\mathbf{A})^{-1}\left(\exp(\Delta\mathbf{A})-\mathbf{I}\right)\Delta\mathbf{B},\end{split}$ (2)
where $\Delta$ is a scalar time step parameter. The discretized state space equations then become:
$\begin{split}h_{t}&=\overline{\mathbf{A}}h_{t-1}+\overline{\mathbf{B}}x_{t},\\ y_{t}&=\mathbf{C}h_{t}.\end{split}$ (3)
This process can be interpreted as generating a structured convolution kernel:
$\begin{split}\overline{\mathbf{K}}&=\left(\mathbf{C}\overline{\mathbf{B}},\ \mathbf{C}\overline{\mathbf{A}}\overline{\mathbf{B}},\ \ldots,\ \mathbf{C}\overline{\mathbf{A}}^{L-1}\overline{\mathbf{B}}\right),\\ \mathbf{y}&=\mathbf{x}*\overline{\mathbf{K}},\end{split}$ (4)
where $L$ is the sequence length, and * denotes 1D convolution.
Selective State Space Models. Traditional SSMs apply fixed parameters $\overline{\mathbf{A}},\overline{\mathbf{B}},\mathbf{C},\Delta$ across all input sequences, which limits their capacity to adapt to diverse and complex input patterns. To address this, Selective SSMs introduce input-dependent dynamics by making parameters such as $\overline{\mathbf{B}},\mathbf{C},\Delta$ functions of the input. This transition from time-invariant to time-variant modeling enhances the flexibility of the architecture, enabling it to selectively focus on salient features in the input sequence.
In our FractalMamba++ architecture, we adopt Selective SSM as the core sequence modeling operator. Its dynamic parameterization allows the model to capture richer temporal dependencies and adapt more effectively to the spatial and contextual variations present in visual inputs.
3.2 Fractal Scanning Curve
Limitations of Linear Curves. As illustrated in Eqn. 1, selecting appropriate input tokens at each time step is critical for effective feature representation modeling in SSMs. Achieving this requires that the serialization of a 2D image into a 1D sequence faithfully retains the image’s inherent structural information. In particular, it is crucial that the serialized patch sequence preserves spatial coherence, so that patches adjacent in the image remain close in the 1D token stream. This proximity facilitates the modeling of both local and global visual patterns. However, existing linear scanning strategies, such as Z-order and Zigzag curves, suffer from fundamental limitations in this regard. While these methods preserve adjacency within individual rows, they often disrupt inter-row spatial relationships. This results in the breakdown of important structural links necessary for capturing the broader spatial layout of the image. Consequently, such linear serialization distorts the spatial structure, impairs semantic consistency, and hinders the model’s ability to generalize across resolutions.
Fractal Curves. To address above limitations, we propose a fractal-based scanning strategy that better preserves the spatial structure of the image across scales. In particular, we adopt the Hilbert curve, a canonical fractal curve constructed via recursive subdivision. Owing to its self-similar and locality-preserving properties, the Hilbert curve ensures that spatially adjacent patches in the 2D image remain close in the 1D sequence, even at different image resolutions. The Hilbert curve begins at a designated point and recursively traverses midpoints along directional vectors $\vec{x}$ and $\vec{y}$, systematically covering the image space. This recursive traversal inherently maintains local continuity, which is essential for accurate representation and analysis. By preserving spatial locality and structural coherence, the fractal scanning mechanism serves as a robust foundation for resolution-scalable visual modeling.
Figure 6: Visualization of correlation between the final output state and intermediate hidden states in the SSM. The leftmost column shows the baseline without CSR, where correlations are concentrated near the sequence end, indicating information fading from earlier patches. The middle and right columns progressively add CSR links (using skip connections). As more CSR connections are introduced, long-range correlations are recovered, demonstrating that CSR effectively alleviates information loss and enhances global context propagation across the serialized patch sequence.
Quantitative Evaluation. To quantitatively evaluate the structural fidelity of different scanning strategies, we introduce a metric called the Structure Distortion Score (SDS). For a given center patch indexed by $i$ (the red one in Fig. 5) in the serialized sequence, we identify its two preceding and two succeeding neighbors, denoted as $\mathcal{N}_{i}=\{i-2,i-1,i+1,i+2\}$. Let $p_{i}\in\mathbb{R}^{2}$ denote the 2D spatial coordinate of patch $i$ in the original image grid. The SDS is defined as the sum of Euclidean distances between the center patch and its four immediate neighbors:
$\text{SDS}(i)=\sum_{j\in\mathcal{N}_{i}}\|p_{i}-p_{j}\|_{2}.$ (5)
A lower SDS indicates stronger preservation of local spatial structure in the serialized sequence.
In Fig. 5, we visualize the SDS value of each patch position on a selected image, with darker colors representing higher distortion levels. It is evident that linear scanning curves suffer from substantial structural distortion in various regions of the image, while the Hilbert fractal curve maintains lower distortion consistently across the grid. This indicates that the fractal curve better preserves spatial adjacency throughout the scanning process. In Tab. I, we report the quantitative SDS statistics across all patches. Notably, 100% of the patches serialized via the Hilbert curve have an SDS below 1.5, while the linear scanning curve fails to meet this threshold for a substantial portion of the sequence. These results further confirm the superior structure-preserving property of fractal-based scanning, which is critical for downstream tasks that require strong spatial consistency, especially under varying image resolutions.
Algorithm 1 Bidirectional Dynamic Programming Process
Input: Input features $\{x_{i}\}_{i=1}^{n}$; State matrix $\{\overline{A}_{i}\}_{i=1}^{n}$; Input matrix $\{\overline{B}_{i}\}_{i=1}^{n}$; Output matrix $\{C_{i}\}_{i=1}^{n}$
1:  Initialization: $\{\mathcal{M}_{i}\}_{i=1}^{n}=\{x_{i}\}_{i=1}^{n}$
2:  # Forward Pass
[…]
4:     if $i==\text{leaf}$ then
5:        $\mathcal{F}_{i}=\overline{B}_{i}x_{i}$
6:     else
7:        $\mathcal{F}_{i}=\overline{B}_{i}x_{i}+\mathcal{F}_{i-1}\overline{A}_{i}$
8:     end if
9:     for $j\in\mathcal{V}_{i}$ (skip nodes) do
10:        $\mathcal{F}_{i}+=\mathcal{F}_{j}\overline{A}_{i}$
11:     end for
12:  end for
13:  # Backward Pass
[…]
15:     if $i==\text{root}$ then
16:        $\mathcal{B}_{i}=\mathcal{F}_{i}$
17:     else
18:        $\mathcal{B}_{i}=\mathcal{B}_{i-1}\overline{A}_{i-1}$
19:     end if
20:     for $j\in\mathcal{U}_{i}$ (skip nodes) do
21:        $\mathcal{B}_{i}+=\mathcal{B}_{j}\overline{A}_{j}$
22:     end for
23:     $y_{i}=C_{i}(\mathcal{F}_{i}+\mathcal{B}_{i})$
24:  end for
Return: Output features $\{y_{i}\}_{i=1}^{N}$
3.3 Cross-State Routing
While fractal scanning curves such as the Hilbert curve significantly improve spatial locality during image serialization, the resulting sequence is still processed uni-directionally in most state space models. As a result, information from early patches may fade as the sequence length increases, particularly in high-resolution settings. This phenomenon limits the model’s ability to retain long-range dependencies and capture global context, which are critical for accurate visual representation. To address this, we propose a skip-connection based mechanism termed Cross-State Routing (CSR). Rather than modifying the scanning order, CSR selectively augments the state transition pathway with a small number of carefully chosen routing links that directly connect distant but semantically related image patches.
These skip routes reinforce weakened long-range transitions and enhance information propagation across the entire image. However, indiscriminately adding skip connections would increase computational complexity and may introduce noise. Therefore, it is crucial to identify a minimal yet highly effective set of skip connections that maximally reinforce critical long-range pathways. To this end, we develop a resolution-agnostic greedy construction strategy that selects structurally beneficial connections while maintaining computational efficiency.
Greedy Construction of Skip Connections. We propose a low-cost greedy strategy to select skip connections that yield the highest structural benefit. Importantly, our method is designed to be independent of image resolution, ensuring scalability and applicability across inputs of varying sizes.
For each patch $p$, we define its Information Aggregation Score $\mathcal{S}(p)$ within a $3\times 3$ window $\mathcal{W}$ as:
$\mathcal{S}(p)=\sum_{p^{’}\in\mathcal{W}}\frac{\mathcal{D}_{f}(p,p^{’})}{\mathcal{D}_{e}(p,p^{’})},$ (6)
where $\mathcal{D}_{f}(\cdot)$ denotes the distance along the fractal curve, and $\mathcal{D}_{e}(\cdot)$ is the 2D Euclidean distance. A lower score indicates better preservation of local structural continuity.
We greedily select the patch with the lowest score as the source node $u$, and then identify a target node $v$ within a fixed window that exhibits the greatest misalignment:
$v=argmin_{p^{’}\in\mathcal{W}}\{\frac{\mathcal{D}_{f}(p,p^{’})}{\mathcal{D}_{e}(p,p^{’})}\}.$ (7)
A skip connection is added between $(u,v)$, and the routing distance is updated as:
$\mathcal{D}_{f}(x,y)=\min\{\mathcal{D}_{f}(x,y),\ \mathcal{D}_{f}(x,u)+\mathcal{D}_{f}(y,v)\}.$ (8)
This process is repeated for $\log(N)$ iterations, where $N$ is the number of patches. As illustrated in Fig. 6, even a small number of skip connections substantially enhances the structural connectivity of the serialized sequence. Our method achieves resolution-agnostic behavior by relying exclusively on relative patch positions within fixed-size local windows, rather than absolute grid dimensions. This design ensures that the skip connection strategy generalizes consistently across resolutions.
Bidirectional Dynamic Programming Process. To incorporate skip connections efficiently, we formulate the Mamba output at position $t$ as follows:
$y_{t}=\mathbf{C}_{t}\sum_{i=1}^{N}x_{i}\overline{\mathbf{B}}_{i}\prod_{j\in\{i\rightarrow t\}}\overline{\mathbf{A}}_{j},$ (9)
where $\{i\rightarrow t\}$ denotes the index set of all vertices along the shortest routing path from patch $i$ to $t$ on the fractal map.
Algorithm 2 Gradient Backpropagation
Input: Gradients $\left\{\frac{\partial loss}{\partial y_{i}}\right\}_{i=1}^{N}$
1:  Initialize: $\mathcal{G}_{i}=\frac{\partial loss}{\partial y_{i}}C_{i}$
[…]
3:     Accumulate forward gradient $\mathcal{G}_{i}$ and propagate via skip nodes
4:  end for
[…]
6:     Accumulate backward path for $\mathcal{M}_{i}$, including skip node contributions
7:     Compute:
  • $\frac{\partial loss}{\partial\overline{B}_{i}}=\mathcal{G}_{i}x_{i}$
  • $\frac{\partial loss}{\partial x_{i}}=\mathcal{G}_{i}\overline{B}_{i}$
  • $\frac{\partial loss}{\partial\overline{A}_{i}}$ as per Eqn. above
  • $\frac{\partial loss}{\partial C_{i}}=\mathcal{F}_{i}+\mathcal{B}_{i}$
8:  end for
Return: Gradients for $\{x_{i},\overline{A}_{i},\overline{B}_{i},C_{i}\}_{i=1}^{N}$
Directly incorporating these skip connections into the standard sequential computation of SSMs would result in a computational burden, as each new connection alters the effective receptive field and routing paths. A naive implementation would require recomputing each patch’s output by enumerating all valid paths, resulting in quadratic complexity. To address this, we reformulate the computation into an efficient process that aggregates contributions from both directions along the curve, while simultaneously leveraging the skip connections.
To reduce computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N+\log N)$, we design a Bidirectional Dynamic Programming Process (BDPP), consisting of a Forward Pass (FP) and Backward Pass (BP), inspired by divide-and-conquer principles. We iteratively compute the forward contributions $\mathcal{F}$ and backward contributions $\mathcal{B}$, with CSR edges dynamically fused during both phases. Final output $y_{t}$ is computed via a gated combination:
$y_{t}=\mathbf{C}_{t}(\mathcal{F}_{t}+\mathcal{B}_{t}).$ (10)
The complete computation flow is presented in Algorithm 1 (inference and forward propagation).
Gradient Backpropagation. We derive gradients of the loss function with respect to inputs $x_{i}$ and state matrices based on Equation 9:
$\frac{\partial y_{t}}{\partial x_{i}}$ $=C_{t}\overline{B}_{i}\prod_{k\in\{i\rightarrow t\}}\overline{A}_{k},$ (11)
$\frac{\partial y_{t}}{\partial\overline{B}_{i}}$ $=C_{t}x_{i}\prod_{k\in\{i\rightarrow t\}}\overline{A}_{k},$ (12)
$\frac{\partial y_{t}}{\partial\overline{A}_{i}}$ $=C_{t}\sum_{k\in\{leaf\rightarrow i\}^{\prime}}x_{k}\overline{B}_{k}\prod_{z\in\{k\rightarrow\hat{i}\}}\overline{A}_{z}\prod_{z\in\{\tilde{i}\rightarrow t\}}\overline{A}_{z},$ (13)
where $\hat{i}$ and $\tilde{i}$ denote the indices before and after $i$, and $\{leaf\rightarrow i\}^{\prime}$ includes all vertices on the path from a leaf to $i$ not passing through $t$. We then compute:
$\frac{\partial loss}{\partial x_{i}}$ $=\sum_{t=1}^{N}\frac{\partial loss}{\partial y_{t}}\cdot\frac{\partial y_{t}}{\partial x_{i}},$ (14)
$\frac{\partial loss}{\partial\overline{B}_{i}}$ $=\sum_{t=1}^{N}\frac{\partial loss}{\partial y_{t}}\cdot\frac{\partial y_{t}}{\partial\overline{B}_{i}},$ (15)
$\frac{\partial loss}{\partial\overline{A}_{i}}$ $=\sum_{t=1}^{N}\frac{\partial loss}{\partial y_{t}}\cdot\frac{\partial y_{t}}{\partial\overline{A}_{i}}.$ (16)
Algorithm 2 shows the flow of gradient backpropagation.
Overall, the CSR mechanism complements fractal serialization by reinforcing information propagation across distant spatial locations. It enables long-range dependency modeling, mitigates early information loss, and ensures efficient computation with only $\mathcal{O}(N+\log N)$ complexity.
TABLE II: Image classification performance (Top-1 Accuracy) on ImageNet-1k under varying input resolutions. FLOPs are measured at input resolution of $224\times 224$.
Method Publication Param. FLOPs Input Resolution
2242 2562 3842 5122 6402 7682 10242 12802 14082 15362
VMamba NeurIPS’24 31M 4.9G 82.5 82.5 82.5 81.1 79.3 76.1 62.3 50.2 45.1 40.9
GrootV NeurIPS’24 30M 4.8G 83.4 83.9 83.6 82.0 80.1 77.6 67.9 52.4 45.0 39.1
MILA NeurIPS’24 25M 4.2G 83.5 83.9 83.5 81.7 79.6 76.8 63.7 49.6 42.8 36.8
MSVMamba NeurIPS’24 33M 4.6G 82.8 82.5 82.3 80.9 78.8 75.1 63.0 54.9 49.6 44.0
Spatial Mamba ICLR’25 27M 4.5G 83.5 83.6 83.0 80.2 77.4 74.4 66.1 53.7 46.4 38.7
Mamba® CVPR’25 29M 4.6G 81.1 45.7 25.4 12.8 7.8 5.3 2.8 1.8 1.6 1.4
MambaVision CVPR’25 32M 4.4G 82.3 81.7 79.8 77.6 74.8 71.2 59.6 46.4 39.7 34.5
FractalMamba AAAI’25 31M 4.8G 83.0 83.5 83.9 83.0 81.8 80.3 76.3 65.9 58.8 52.1
FractalMamba++ Year’25 30M 4.8G 83.0 83.3 84.1 83.9 83.0 81.9 78.8 74.3 71.3 67.5
MSVMamba NeurIPS’24 12M 1.5G 79.8 80.1 80.0 78.3 75.8 72.0 59.4 43.9 36.5 29.9
Efficient VMamba AAAI’25 11M 1.3G 78.7 79.6 79.5 77.3 75.2 72.4 64.2 54.1 42.6 38.3
FractalMamba++ Year’25 11M 1.6G 79.5 80.6 82.0 81.3 80.1 78.3 73.3 66.3 61.7 56.1
MSVMamba NeurIPS’24 7M 0.9G 77.3 77.7 77.4 75.0 71.7 65.8 48.0 31.0 23.8 18.3
ViM ICML’24 7M 1.5G 76.1 76.3 70.4 67.4 51.4 30.6 16.1 7.2 4.1 1.8
Efficient VMamba AAAI’25 6M 0.8G 76.5 76.9 76.5 73.8 70.4 65.8 52.0 36.2 29.4 24.1
FractalMamba++ Year’25 7M 1.0G 77.3 78.4 79.5 78.4 76.4 73.7 66.5 55.2 48.1 42.5
3.4 Positional-Relation Capture
While fractal scanning curves such as the Hilbert curve effectively preserve spatial locality during the serialization of 2D image patches into 1D sequences, they still present certain limitations. In particular, structural continuity can be disrupted at inflection points of the curve, such as corners, where spatially adjacent patches may be serialized into distant positions. This misalignment impairs the model’s capacity to fully capture local context, especially under varying image resolutions. To address this issue, we incorporate a complementary positional encoding strategy based on Rotary Positional Embedding (RoPE) [49], which enhances the model’s sensitivity to relative positional relationships within the serialized sequence. RoPE is a recent advancement in relative positional encoding that addresses the limitations of conventional Relative Position Bias (RPB) [50]. In RPB, position-dependent biases are introduced after computing the query-key similarity, which restricts their influence on the similarity computation itself. RoPE overcomes this limitation by injecting positional information directly into the similarity function through complex rotations applied to the query and key vectors.
Given the $n$-th query $\mathbf{q}_{n}\in\mathbb{R}^{1\times d_{\text{head}}}$ and the $m$-th key $\mathbf{k}_{m}\in\mathbb{R}^{1\times d_{\text{head}}}$, RoPE applies a rotation in the complex plane:
$\mathbf{q}_{n}^{\prime}=\mathbf{q}_{n}e^{in\theta},\quad\mathbf{k}_{m}^{\prime}=\mathbf{k}_{m}e^{im\theta},$ (17)
and computes the attention score as:
$\mathbf{A}_{(n,m)}^{\prime}=\operatorname{Re}\left[\mathbf{q}_{n}\mathbf{k}_{m}^{*}e^{i(n-m)\theta}\right],$ (18)
where $\operatorname{Re}[\cdot]$ denotes the real part and is the complex conjugate. This formulation introduces a rotation-phase difference proportional to the relative distance $(n-m)$, thereby encoding relative positions directly into token similarity.
To implement this in practice, the query and key vectors are reinterpreted as complex pairs: every $(2t)$-th and $(2t+1)$-th dimensions are treated as real and imaginary components. RoPE further defines multiple rotation frequencies across channels using:
$\theta_{t}=10000^{-t/(d_{\text{head}}/2)},$ (19)
which yields a rotation matrix:
$\mathbf{R}(n,t)=e^{i\theta_{t}n},$ (20)
applied via element-wise Hadamard product:
$\overline{\mathbf{q}}_{n}^{\prime}=\overline{\mathbf{q}}_{n}\circ\mathbf{R},\quad\overline{\mathbf{k}}_{m}^{\prime}=\overline{\mathbf{k}}_{m}\circ\mathbf{R},\quad\mathbf{A}^{\prime}=\operatorname{Re}\left[\overline{\mathbf{q}}_{n}^{\prime}\cdot\overline{\mathbf{k}}_{m}^{\prime*}\right].$ (21)
This rotation mechanism enables continuous and position-aware token interactions, which are especially beneficial in vision models that serialize spatially structured data. In our setting, RoPE acts as a targeted enhancement to the fractal scanning process: by embedding explicit relative position information into the serialized sequence, it compensates for the structural distortions introduced at inflection points of the Hilbert curve. This positional reinforcement strengthens the local consistency of spatially adjacent patches that are otherwise distant in the serialization, thereby mitigating the residual locality breaks of fractal ordering. Therefore, RoPE complements the fractal-based design and contributes to reliable spatial reasoning in Vision Mamba.
TABLE III: Inference speed comparison of Mamba-based models on ImageNet-1K at $224\times 224$ input. All timings are measured on a single H800 GPU with batch size 32.
Method Param. FPS
VMamba 31M 1916
GrootV 30M 315
MILA 25M 47
MSVMamba 33M 684
Spatial Mamba 27M 957
Mamba® 29M 1044
MambaVision 32M 3426
FractalMamba++ 30M 302
MSVMamba 12M 1000
Efficient VMamba 11M 2187
FractalMamba++ 11M 511
MSVMamba 7M 1181
ViM 7M 398
Efficient VMamba 6M 2380
FractalMamba++ 7M 613
4 Experiments
We conduct a series of experiments to evaluate and compare FractalMamba++ against several established Mamba-based benchmark models across a diverse range of vision tasks. These tasks include image classification, object detection, remote sensing binary change object detection, and semantic segmentation. Each task requires a different input resolution, which provides a comprehensive testbed to evaluate the multi-scale adaptability of our model. In particular, the remote sensing binary change object detection task involves high-resolution input images of size 1024 $\times$ 1024. This setting places strong demands on the model’s ability to maintain spatial consistency and capture long-range structural relationships. All experiments are conducted using 8 NVIDIA H800 GPUs.
4.1 Image Classification
We evaluate the classification performance of FractalMamba++ on the ImageNet-1K dataset [51], comparing it against a diverse set of representative Mamba-based vision backbones. To assess the scalability and robustness of our model under different image scales, we conduct testing across a broad range of input resolutions, from $224^{2}$ to $1536^{2}$, covering both standard and extreme high-resolution scenarios. FractalMamba++ models are trained from scratch for 300 epochs with a 20-epoch warm-up period. The training uses a batch size of 1024 and the AdamW optimizer [52], with $\beta_{1}=0.9$, $\beta_{2}=0.999$, momentum of 0.9, a cosine decay learning rate schedule starting at $1\times 10^{-3}$, and a weight decay of 0.05. We also apply label smoothing with a factor of 0.1 and exponential moving average (EMA) to stabilize training.
We benchmark FractalMamba++ at three model sizes: 30M, 11M, and 7M parameters, enabling fair comparisons under different model capacities. For the 30M-scale models, we compare with VMamba [26], GrootV [27], MILA [33], MSVMamba [34], SpatialMamba [35], MambaVision [36], Mamba® [37], and FractalMamba [32]. At the 11M and 7M scales, we evaluate against Efficient VMamba [38], MSVMamba [34], and ViM [25]. The detailed performance comparisons are reported in Table II. All evaluations are conducted under a unified testing environment using official checkpoints from the respective repositories to ensure fairness and reproducibility. FLOPs are also reported to reflect computational efficiency, providing a comprehensive view of accuracy–efficiency trade-offs.
FractalMamba++ consistently achieves competitive or superior performance across all tested resolutions. At lower resolutions (e.g., $224^{2}$, $384^{2}$), it performs on par with state-of-the-art Mamba-based models. However, as input resolution increases, FractalMamba++ shows significantly slower degradation in accuracy compared to all baselines. For instance, at $1024^{2}$, the 30M variant of FractalMamba++ achieves 78.8% top-1 accuracy, outperforming SpatialMamba (66.1%), GrootV (67.9%), and the original VMamba (62.3%). Similar trends are observed at 11M and 7M scales. This superior scalability highlights the effectiveness of our design. The fractal-based patch serialization preserves spatial locality across scales, while the CSR and PRC modules enhance long-range context propagation and mitigate structural disruptions caused by curve inflections. Together, these modules allow FractalMamba++ to maintain strong performance under extreme-resolution inputs, confirming its multi-scale adaptability and robustness for real-world visual applications.
To complement accuracy evaluations, we further compare the inference efficiency of various Mamba-based models on ImageNet-1K at $224\times 224$ resolution. As shown in Table III, we report frames per second (FPS) on a single H800 GPU with batch size 32. FractalMamba++ achieves competitive inference speeds within each parameter scale, while offering consistently superior multi-resolution performance. In the $\sim$30M group, FractalMamba++ runs at 302 FPS, slower than VMamba (1916 FPS) and MambaVision (3426 FPS), but significantly faster than MILA (47 FPS) and on par with GrootV (315 FPS). In the medium-scale group ($\sim$11M), FractalMamba++ achieves 511 FPS, outperforming ViM (398 FPS) in the small-scale group ($\sim$7M), while Efficient VMamba maintains the highest FPS across all groups. Although FractalMamba++ incurs moderate latency due to fractal serialization and additional modules (CSR, PRC), it still delivers a favorable trade-off between speed and accuracy, especially under high-resolution scenarios where other models degrade sharply. Notably, the current implementation of the CSR module involves custom CUDA operators, which, while functional, have not been fully optimized. This suggests that there remains considerable headroom for further acceleration through kernel-level refinement. This highlights the practical deployment potential of FractalMamba++ in resource-constrained settings.
TABLE IV: Object detection and instance segmentation performance on COCO using Mask R-CNN under 1× and 3× training schedules. FLOPs are measured at input resolution of $1280\times 800$.
Method Param. FLOPs Mask R-CNN 1× Schedule Mask R-CNN 3× Schedule
$AP^{\text{box}}$ $AP^{\text{box}}_{50}$ $AP^{\text{box}}_{75}$ $AP^{\text{mask}}$ $AP^{\text{mask}}_{50}$ $AP^{\text{mask}}_{75}$ $AP^{\text{box}}$ $AP^{\text{box}}_{50}$ $AP^{\text{box}}_{75}$ $AP^{\text{mask}}$ $AP^{\text{mask}}_{50}$ $AP^{\text{mask}}_{75}$
VMamba 42M 286G 46.5 68.5 50.7 42.1 65.5 45.3 48.5 69.9 52.9 43.2 66.8 46.3
GrootV * 265G 47.0 69.4 51.5 42.7 66.4 46.0 49.0 70.8 54.0 43.8 67.6 47.1
MSVMamba 53M 252G 46.9 68.8 51.4 42.2 65.6 45.4 48.3 69.5 53.0 43.2 66.8 46.9
SpatialMamba 46M 261G 47.6 69.6 52.3 42.9 66.5 46.2 49.3 70.7 54.3 43.6 67.6 46.9
DefMamba * 268G 47.5 69.6 51.7 42.8 66.3 46.2 * * * * * *
FractalMamba 41M 266G 46.8 68.7 50.8 42.4 65.9 45.8 48.5 70.0 53.2 43.3 67.1 46.2
FractalMamba++ 41M 260G 47.8 69.8 52.5 43.2 66.8 46.6 49.5 71.0 54.6 44.1 67.9 47.5
MSVMamba 32M 201G 43.8 65.8 47.7 39.9 62.9 42.9 46.3 68.1 50.8 41.8 65.1 44.9
Efficient VMamba 31M 197G 39.3 61.8 42.8 36.7 58.9 39.2 41.6 63.9 45.6 38.2 60.8 40.7
FractalMamba++ 31M 199G 44.1 66.2 48.0 40.3 63.2 43.4 46.8 68.5 51.3 42.2 65.4 45.3
4.2 Object Detection
We evaluate FractalMamba++ on object detection and instance segmentation using the MS COCO2017 dataset [53]. Experiments are conducted using the MMDetection toolbox [54], following the standard training setup established by Swin-T [28] with the Mask R-CNN detector [55]. All models are initialized from ImageNet-1K pre-trained weights and fine-tuned under both the 1× (12 epochs) and 3× (36 epochs) schedules. Optimization uses AdamW [52] with an initial learning rate of $1\times 10^{-4}$, which is reduced at the 9th and 11th epochs in the 1× setting. A drop path rate of 0.2 is used, and multi-scale training along with random horizontal flipping is applied. Input resolution is fixed at $1280\times 800$ with a batch size of 16.
Table IV presents detection and segmentation results for FractalMamba++ and several representative Mamba-based models, including VMamba [26], MSVMamba [34], Efficient VMamba [38], SpatialMamba [35], DefMamba [56], and FractalMamba [32]. Performance is reported in terms of mean average precision (mAP) for bounding boxes and instance masks, including scores at 50 and 75 percent IoU thresholds. FractalMamba++ achieves consistent improvements under both training settings. Under the 1× schedule, the 41M model obtains 47.8 mAP for box prediction and 43.2 mAP for mask prediction, outperforming Spatial Mamba (47.6 and 42.9), MSVMamba (46.9 and 42.2), and VMamba (46.5 and 42.1). Under the 3× schedule, FractalMamba++ further improves to 49.5 for box mAP and 44.1 for mask mAP, achieving the highest performance among comparable models. Similar trends are observed at the 31M scale, where FractalMamba++ achieves 46.8 for box mAP and 42.2 for mask mAP, compared with 46.3 and 41.8 from MSVMamba. These results confirm the effectiveness and generalizability of FractalMamba++ across dense prediction tasks. The fractal-based serialization improves spatial structure retention, while the CSR and PRC modules enable stronger context modeling and structural continuity, which are essential for high-resolution visual understanding.
TABLE V: Semantic segmentation results on ADE20K using UPerNet with input resolution $512\times 2048$. FLOPs are computed under the same setting. SS and MS denote single-scale and multi-scale inference.
Method Param. FLOPs mIoU (SS) mIoU (MS)
VMamba 55M 946G 47.3 48.3
GrootV * 941G 48.5 49.4
MSVMamba 65M 942G 47.6 48.5
Spatial Mamba 57M 936G 48.6 49.4
DefMamba 65M 946G 48.7 49.6
MambaVision 55M 945G 46.0 *
Mamba® 56M * 45.3 *
FractalMamba 53M 942G 48.0 48.9
FractalMamba++ 62M 957G 48.8 49.7
MSVMamba 42M 875G 45.1 45.4
ViM 46M * 44.9 *
FractalMamba++ 44M 889G 45.8 46.2
Efficient VMamba 29M 505G 41.5 42.1
Local-ViM 36M 181G 43.4 44.4
PlainMamba 35M 174G 44.1 44.6
FractalMamba++ 37M 860G 44.5 45.0
4.3 Semantic Segmentation
We assess the segmentation performance of FractalMamba++ on the ADE20K dataset [57] using the UPerNet framework [58], following the training setup established by Swin [28]. The backbone is initialized with ImageNet-1K pre-trained weights and fine-tuned for 160,000 iterations using the AdamW optimizer [52] with a learning rate of $6\times 10^{-5}$ and a batch size of 16. The training input resolution is $512\times 512$, while FLOPs are measured at $512\times 2048$ for consistency with prior work. We report both single-scale (SS) and multi-scale (MS) inference results. Additionally, we include $640\times 640$ inference to evaluate the model’s robustness under larger input scales.
Table V reports results across three model capacities. At the Tiny scale (62M parameters), FractalMamba++ achieves 48.8 mIoU (SS) and 49.7 mIoU (MS), outperforming prior models such as Spatial Mamba (48.6/49.4), MSVMamba (47.6/48.5), and FractalMamba (48.0/48.9). At the Medium scale (44M), it achieves 45.8/46.2, exceeding MSVMamba-M (45.1/45.4) and Efficient VMamba (41.5/42.1). At the Nano scale (37M), it obtains 44.5/45.0, outperforming Local-ViM (43.4/44.4) and PlainMamba (44.1/–). These results demonstrate the strong segmentation capability of FractalMamba++ across different model sizes. The consistent performance gain over existing Mamba-based backbones highlights the benefits of its fractal-based serialization, long-range routing via CSR, and enhanced positional encoding through PRC.
TABLE VI: Performance and efficiency comparison on binary change detection using the LEVIR-CD+ dataset (input resolution: 1024×1024).
Method Param. FLOPs IoU Precision Recall KC F1
ChangeMamba 17M 46G 78.6 88.8 87.3 87.5 88.0
ChangeMamba-S 50M 115G 78.3 89.2 86.5 87.3 87.8
ChangeMamba-B 85M 179G 79.2 89.2 87.6 87.9 88.4
FractalMamba 35M 55G 80.0 89.3 88.4 88.4 89.9
FractalMamba++ 35M 56G 80.5 90.0 88.5 88.7 90.0
4.4 Remote Sensing Binary Change Detection
We evaluate the effectiveness of FractalMamba++ on the binary change detection (BCD) task using the LEVIR-CD+ dataset, following the standard protocol introduced in ChangeMamba [59]. LEVIR-CD+ is an enhanced version of the original LEVIR-CD dataset and consists of 985 pairs of very high-resolution aerial image patches at a spatial resolution of 0.5 meters per pixel, with each image sized at $1024\times 1024$ pixels.
To ensure a fair comparison, we adopt the same Change-Decoder as used in ChangeMamba and initialize the backbone with ImageNet-pretrained weights. The training process uses the AdamW optimizer with a learning rate of $1\times 10^{-4}$, weight decay of 0.005, and a batch size of 16. The model is evaluated using five widely adopted metrics: Intersection over Union (IoU), Precision, Recall, Kappa Coefficient (KC), and F1 Score. All metrics are expressed as percentages, where higher values indicate better performance. As shown in Table VI, FractalMamba++ achieves the best performance across all evaluation metrics. Specifically, it attains an IoU of 80.5 and an F1 Score of 90.0, outperforming its predecessor FractalMamba (80.0 IoU / 89.9 F1) despite using a similar number of parameters and computational cost (35M / 56G). Compared to ChangeMamba-B, which has 85M parameters and 179G FLOPs, FractalMamba++ demonstrates superior accuracy with significantly lower complexity. These results highlight the efficiency and representational strength of FractalMamba++ in high-resolution remote sensing change detection tasks.
4.5 Ablation Studies
We conduct ablation studies of our FractalMamba++ across various resolutions on the ImageNet-1K dataset. Classification performance is evaluated under multiple input resolutions: 2242, 3842, 6402, and 10242.
TABLE VII: Ablation study on the effectiveness of proposed modules. VMamba is used as the baseline. Each component is incrementally added, and classification performance is reported across input resolutions.
Fractal CSR PRC Model 2242 3842 6402 10242
VMamba 82.5 82.5 79.3 62.9
+Fractal 82.7 82.8 80.6 72.3
+CSR 83.0 83.7 82.1 76.2
+PRC 83.0 84.1 83.0 78.8
4.5.1 Effectiveness of Proposed Modules
We conduct an ablation study to assess the individual contributions of the three key components in FractalMamba++: Fractal-based patch serialization (Fractal), Cross-State Routing (CSR), and Positional-Relation Capture (PRC), using VMamba as the baseline model. These components are incrementally integrated, and classification performance is evaluated at multiple resolutions ranging from $224^{2}$ to $1024^{2}$.
As shown in Table VII, introducing the Fractal curve alone yields consistent improvements, particularly under higher resolutions (e.g., +9.4% at $1024^{2}$), indicating its effectiveness in maintaining spatial structure during serialization. Adding the CSR mechanism further enhances performance, especially at larger scales (e.g., +3.9% at $1024^{2}$), by facilitating long-range interactions and mitigating information fading. Incorporating the PRC module provides the final boost, bringing performance to 84.1% at $224^{2}$ and 78.8% at $1024^{2}$, highlighting its ability to reinforce relative positional cues and address residual structural disruptions caused by curve inflections. These results validate the complementary nature of the proposed modules and demonstrate their effectiveness in enhancing the resolution adaptability and spatial modeling capacity of FractalMamba++.
TABLE VIII: Ablation on the effectiveness of fractal serialization in improving the multi-resolution classification performance of ViM and VMamba. All models are evaluated on ImageNet-1K with input resolutions from 2242 to 10242.
Backbone
Fractal
Curve
2242 3842 6402 10242
ViM 76.1 70.4 51.4 16.1
ViM 77.9 75.2 62.3 39.7
VMamba 82.5 82.5 79.3 62.9
VMamba 82.7 82.8 80.6 72.3
4.5.2 Effectiveness of Fractal Curves in Mamba Backbone
To evaluate the general applicability of fractal-based patch serialization, we apply the Hilbert fractal curve to two representative vision architectures: ViM and VMamba. For each backbone, we compare the standard version (without fractal serialization) with its fractal-augmented variant. All models are trained and evaluated on ImageNet-1K under multiple input resolutions ranging from $224^{2}$ to $1024^{2}$.
Table VIII shows that fractal serialization significantly improves classification performance, particularly under high-resolution inputs. For ViM, the top-1 accuracy at $1024^{2}$ increases from 16.1% to 39.7%, indicating that the fractal curve effectively preserves spatial continuity that would otherwise be lost in naive raster scanning. Similarly, for VMamba, accuracy improves from 62.9% to 72.3% at $1024^{2}$. These results confirm that fractal-based patch serialization enhances the multi-resolution robustness of both convolution-free and SSM-based architectures, demonstrating its generalizability and effectiveness across different model families.
TABLE IX: Ablation study on adding CSR to ViM and VMamba. Evaluation is conducted on ImageNet-1K across input resolutions.
Backbone CSR 2242 3842 6402 10242
ViM 76.1 70.4 51.4 16.1
ViM 76.5 71.2 53.0 18.4
VMamba 82.5 82.5 79.3 62.9
VMamba 82.6 82.8 80.1 64.3
4.5.3 Effectiveness of CSR in Mamba Backbone
To further assess the general utility of the proposed CSR mechanism, we integrate it into two baseline architectures: ViM and VMamba. As shown in Table IX, adding CSR results in minor accuracy improvements at low resolutions and moderate gains at higher resolutions. Specifically, VMamba benefits more from CSR at 6402 and 10242, where global dependency modeling becomes increasingly important. However, the overall performance boost remains limited when used alone, highlighting the need for joint design with spatial-aware serialization methods like fractal curves.
TABLE X: Ablation study on different types of fractal curves in VMamba. Hilbert, Coil, and Meurthe curves are evaluated under varying input resolutions.
Fractal Curve 2242 3842 6402 10242
Hilbert 83.0 83.3 83.0 78.8
Coil 82.8 83.1 82.6 78.5
Meurthe 82.9 83.0 82.8 78.5
4.5.4 Comparison of Fractal Curve Types
To assess the impact of different fractal serialization strategies, we replace the Hilbert curve in VMamba with two alternative fractal curves: Coil and Meurthe. As shown in Table X, Hilbert consistently outperforms the other two curves across all resolutions. At 10242, Hilbert achieves 78.8 accuracy, while Coil and Meurthe reach 78.3 and 78.1, respectively. Although the performance gap is small, the results suggest that Hilbert is slightly more effective in preserving spatial locality during serialization. Nonetheless, the competitive results of Coil and Meurthe demonstrate that our framework is robust to the choice of fractal curve.
TABLE XI: Comparison between fractal-based modeling and multi-scale training. All models are trained on ImageNet-1K using VMamba as the backbone. For MS-Train, we randomly sample resolutions from {224, 384, 512} during training.
Method 2242 3842 6402 10242
MS-Train 83.1 83.2 80.4 72.0
Fractal Only 83.0 83.3 83.0 78.8
4.5.5 Fractal-based Modeling versus Multi-scale Training
To evaluate whether fractal-based modeling can serve as a substitute for multi-scale training, we conduct experiments on ImageNet-1K using GrootV as the backbone. In the multi-scale training (MS-Train) setup, input resolutions are randomly sampled from ${224^{2},384^{2},512^{2}}$ during training. In contrast, the fractal-only setup trains on a fixed resolution of $224^{2}$ but applies fractal serialization to preserve spatial structure.
Table XI shows that both methods perform similarly at low resolutions. For example, at $224^{2}$ and $384^{2}$, the accuracy of fractal modeling is comparable to that of MS-Train. However, as input resolution increases, fractal-based modeling demonstrates significantly better generalization, achieving 83.0% at $640^{2}$ and 78.8% at $1024^{2}$, outperforming MS-Train by 2.6 and 6.8 points, respectively. These results suggest that fractal serialization can better maintain structural consistency across scales, offering a resolution-robust alternative to traditional multi-scale training.
5 Conclusion
This paper presents FractalMamba++, a novel vision backbone that enhances the Mamba architecture with improved adaptability to spatial structure and resolution variability. Motivated by the limitations of existing linear serialization strategies, we introduce a fractal-based patch scanning approach that better preserves local spatial relationships during 2D-to-1D conversion. To further address long-range dependency degradation and structural discontinuities introduced by fractal inflection points, we incorporate two dedicated modules: Cross-State Routing (CSR) for enhanced global information propagation and Positional-Relation Capture (PRC) for explicit modeling of spatial adjacency. Extensive experiments demonstrate the effectiveness and robustness of FractalMamba++ across a wide range of visual tasks and resolutions. Our method consistently outperforms existing Mamba-based backbones in multi-resolution image classification, achieving competitive performance at low resolutions while maintaining strong accuracy under extreme-resolution inputs.